# A Disentangled Representation Learning Framework for Low-altitude Network Coverage Prediction

### v1

@@ Abstract
The expansion of the low-altitude economy has underscored the significance of Low-Altitude Network Coverage (LANC) prediction for designing aerial corridors. While accurate LANC forecasting hinges on the antenna beam patterns of Base Stations (BSs), these patterns are typically proprietary and not readily accessible. Operational parameters of BSs, which inherently contain beam information, offer an opportunity for data-driven low-altitude coverage prediction. However, collecting extensive low-altitude road test data is cost-prohibitive, often yielding only sparse samples per BS. This scarcity results in two primary challenges: imbalanced feature sampling due to limited variability in high-dimensional operational parameters against the backdrop of substantial changes in low-dimensional sampling locations, and diminished generalizability stemming from insufficient data samples. To overcome these obstacles, we introduce a dual strategy comprising expert knowledge-based feature compression and disentangled representation learning. The former reduces feature space complexity by leveraging communications expertise, while the latter enhances model generalizability through the integration of propagation models and distinct subnetworks that capture and aggregate the semantic representations of latent features. Experimental evaluation confirms the efficacy of our framework, yielding a $7\%$ reduction in error compared to the best baseline algorithm. Real-network validations further attest to its reliability, achieving practical prediction accuracy with MAE errors at the $5\ \mathrm{dB}$ level.
@@ I Introduction
With the burgeoning development of the low-altitude economy, researchers have redirected their attention to LANC prediction[2]. In contrast to terrestrial networks, low-altitude networks will be planned in airspace with fewer buildings, leading to the predominance of received signals from direct paths. Consequently, the actual LANC is largely influenced by the multi-beam capabilities of base stations (BSs) and the collaborative effect of multiple BSs[20]. This shifts the focus from conventional environmental scattering to base station antenna configurations and beamforming patterns. Moreover, many terrestrial models treat the vertical dimension as negligible, making them unsuitable for the three-dimensional nature of low-altitude networks. The traditional terrestrial solutions, which are primarily designed for ground-level environments, are inadequate for addressing these challenges. As a result, there is a clear need for novel approaches specifically tailored to the unique requirements of low-altitude networks.
@@
  • Experiments with Real-world Network Data: We validated the proposed framework through real-world low-altitude data collection and network training conducted in Nanchang, Jiangxi Province, China. Ablation studies demonstrate our framework reduces the mean absolute error (MAE) by approximately $20\%$ compared to traditional DNNs. Furthermore, our method outperforms the best baseline algorithm with a $7\%$ reduction in error. In practical LANC prediction of an unknown area, the framework achieves a measured MAE error at the $5\ \mathrm{dB}$ level, demonstrating its reliability and applicability in real-world scenarios.
@@ II-A BS and Its Operational Parameters
In scenarios such as BS pre-planning or where low-altitude sparse data collection is absent, predicting low-altitude coverage based on operational parameters becomes a prevalent approach. For such a BS of interest, the operational parameters crucially linked to coverage can be systematically categorized based on their roles in determining coverage as follows:
@@
For ease of description, the aforementioned categories and the included operational parameters are defined in Table I.
TABLE I: Category of Operational Parameters and Corresponding Mathematical Expression
@@ II-B Low-altitude Coverage for Area of Interest
To further explore the LANC provided by the BS, consider a nearby area with dimensions $U\times W\times V$ m3, corresponding to its longitudinal, latitudinal, and altitudinal extents. This area can be subdivided into $N$ cubes, where $N=N_{U}\times N_{W}\times N_{V}$ and $N_{U}$, $N_{W}$, and $N_{V}$ represent the counts of discretization along the longitude, latitude, and altitude axes, respectively. The set of cubes can thus be defined as $\mathbf{R}=\{1,\cdots,r_{n},\cdots,r_{N}\}$. For cube $r_{n}$, the absolute position vector can be defined as $\mathbf{x^{(R)}_{n}}=[x^{(R)}_{n,1},x^{(R)}_{n,2},x^{(R)}_{n,3}]^{T}$, where $x^{(R)}_{n,1}$, $x^{(R)}_{n,2}$ and $x^{(R)}_{n,3}$ are longitude, latitude, and altitude, respectively. Furthermore, the absolute position array for all cubes is defined as $\mathbf{X^{(R)}}=[\mathbf{x^{(R)}_{1}},\cdots,\mathbf{x^{(R)}_{n}},\cdots,\mathbf{x^{(R)}_{N}}]\in\mathbb{R}^{3\times N}$.
Determined by the AAU type, the BS will emit $M$ beams. For any given cube $n$, $M$ Synchronization-Signaling-Block-RSRP (SSB-RSRP) measurements can be obtained, along with a maximum SSB-RSRP, denoted as SS-RSRP. These measurements can be represented as $\mathbf{p_{n}}=[p_{n,0},p_{n,1},\cdots,p_{n,m},\cdots,p_{n,M}]^{T}\in\mathbb{R}^{M\times 1}$, where $p_{n,0}$ is the SS-RSRP for the $n$-th cube, and $p_{n,m}$ represents the $m$-th SSB-RSRP for the $n$-th cube. Consequently, the coverage for all cubes within the area can be expressed as $\mathbf{P}=[\mathbf{p_{1}},\cdots,\mathbf{p_{n}},\cdots,\mathbf{p_{N}}]\in\mathbb{R}^{M\times N}$.
@@ II-C Problem Formulation
For the whole area, the prediction of $\mathbf{P}$ can be denoted as $\mathbf{\hat{P}}=[\mathbf{\hat{p}_{1}},\cdots,\mathbf{\hat{p}_{n}},\cdots,\mathbf{\hat{p}_{N}}]\in\mathbb{R}^{M\times N}$. Finally, the problem of coverage prediction can be formulated as
$\min_{g}L(\mathbf{P},\mathbf{\hat{P}}),$ (3)
@@ III-A Expert-knowledge-based Feature Compressing
where $\theta_{n}^{(H)}$ and $\theta_{n}^{(V)}$ are the horizontal and vertical angles of cube $r_{n}$ relative to the BS’s true north direction, respectively. The elements $\Delta x^{(R)}_{n,1}$, $\Delta x^{(R)}_{n,2}$, and $\Delta x^{(R)}_{n,3}$ represent the first, second, and third components of $\Delta\mathbf{x^{(R)}_{n}}$, respectively. These components signify the differences in longitude, latitude, and altitude dimensions between cube $r_{n}$ and the considered BS. Eq. (4) returns the horizontal relative angle of $r_{n}$ with respect to the true north direction of the considered BS, with a range of $[-\pi,\pi]$ (in radians).
@@
$D_{n}=\sqrt{{\Delta x^{(R)}_{n,1}}^{2}+{\Delta x^{(R)}_{n,2}}^{2}+{\Delta x^{(R)}_{n,3}}^{2}}$ (8)
@@
Here, $p^{(T)}$ represents the transmission power of the SSB signal, $G^{(T)}_{n,m}$ denotes the directional gain of the BS for the $m$th beam at grid $r_{n}$, and $G^{(R)}$ is the gain of the receiver antenna. $PL_{n,m}$ is the path fading between cube $r_{n}$ and corresponding beam $m$ [^3] In the context of LANC, different beams may follow distinct paths, leading to varied path fading. Given that subsequent experiments are conducted using the same single-antenna UAV, without loss of generality, we simplify the receiver antenna to an omnidirectional antenna.
Theoretically, $p^{(T)}$ can be obtained from $x^{(A)}_{1}-10\log_{10}x^{(A)}_{2}$. However, due to the actual SSB signal bandwidth utilization not being 100% and its susceptibility to the BS’s carrier frequency and bandwidth, represented by $\sigma$, the actual $p^{(T)}$ is given by
@@
By incorporating expert knowledge, the transformation of the prediction task has addressed the issue of imbalanced feature variation. Simultaneously, it significantly reduces the dimensionality of the input variables, thereby ensuring an enhancement in coverage prediction performance. Next, we will introduce the corresponding disentangled representation network design.
@@ IV-A1 Definitions
Figure 4: Full model and prior constrained model.
@@ IV-A2 Decomposing model for comparable learning capability
We try to establish that our proposed prior-constrained model structure can maintain a comparable learning capability to traditional DNNs. From an angle of deep learning practitioners, knowing how $\mathbf{x}$ is independently composed of $[\mathbf{x}^{1T},\dots,\mathbf{x}^{KT}]$, we propose to design dedicated model heads for each of the components as shown in Fig. 4. Then, to make use of the full data, we aggregate the output of the heads as the final feature by concatenation. This procedure is equivalent to a constrained model. To see this, take a toy example of the linear layer in a full model, transforming a vector of $d$ dimensions into one of $d_{1}$ dimensions. The weight of this layer can be viewed as a matrix $\mathbf{w}\in\mathbb{R}^{d\times d_{1}}$. To gain a corresponding constrained model, some dimension of $\mathbf{w}$ can be constrained to $\{0\}$.
@@
$C$ is a binary random variable, takes value $1$ if a $\mathbf{w}\in\mathcal{W}_{c}$ else 00.
@@
TABLE II: Simple Description of Dataset
Area Index Longitude Latitude Length (m) Width (m) Height (m) Total Sample Number Corresponding BS Number
1 115.69 28.61 4006 4273 150,300,500 43919 33
@@
TABLE III: Benchmarks for Ablation Study
@@ V-A Dataset
To evaluate the performance of the proposed method, initial low-altitude road tests were conducted across three distinct areas within Nanchang City, Jiangxi Province. For these tests, a UAV equipped with a road test terminal flew at varying altitudes (150m, 300m, 500m). The terminal was capable of receiving multiple SSB-Beams’ RSRP and PCI values from the downlink of the serving cell BSs. Utilizing the PCI values, the terminals identified and matched the nearest BS with the corresponding PCI within the area. The samples collected from each region during the tests are presented in Table II.
@@ V-C1 Network Structure Comparison
To validate the advanced nature of the proposed network structure, we established three benchmarks listed in Table III for ablation studies. In all benchmarks, the number of neurons in the hidden layers is set to 256, with the ReLU function serving as the activation function. For all benchmarks, the learning rate is set to 0.001, with a maximum of 100 epochs. Early stopping is implemented, terminating training when the loss on either the training or validation set decreases by no more than 1% over ten consecutive epochs. For benchmark 1, the hidden layer counts for distance fading, frequency fading, and antenna gain networks are all 5. In contrast, benchmarks 2 and 3 feature 6 hidden layers. Compared to the benchmark proposed in this paper, benchmark 2 utilizes decoupled features as inputs and employs an MLP to perform predictions as depicted in Eq. (11). Conversely, benchmark 3 employs original features as inputs and uses an MLP to execute predictions according to Eq. (2).
@@
Figure 6: Boxplot of MAE performance for different algorithms across 20 random splits at sampling rates of $[10\%,20\%,30\%,40\%,50\%,60\%,70\%]$, totally 140 MAE data points per algorithm.
Figure 7: Boxplot of MAE performance for different algorithms across 20 random splits at sampling rates of $[1\%,2\%,3\%,4\%,5\%,6\%]$, totally 120 MAE data points per algorithm.
To demonstrate the effectiveness of the disentangled feature and network modules, we initially conducted experiments with 20 different dataset splits at sampling rates of $[10\%,20\%,30\%,40\%,50\%,60\%,70\%]$. For each sampling rate, we plotted the average MAE across benchmarks 1-3, as illustrated in Fig. 5. It is observable that the prediction errors for benchmarks 3, 2, and 1 progressively decrease, which validates the efficacy of the module design proposed in this study. Furthermore, as the sampling rate increases, the trends of the MAE curves for all three methods are consistently aligned, with overall performance gradually improving. This indicates that with an increase in training dataset size, the algorithms can learn a more precise distribution of the original dataset.

### v2

@@ Abstract
The expansion of the low-altitude economy has underscored the significance of Low-Altitude Network Coverage (LANC) prediction for designing aerial corridors. While accurate LANC forecasting hinges on the antenna beam patterns of Base Stations (BSs), these patterns are typically proprietary and not readily accessible. Operational parameters of BSs, which inherently contain beam information, offer an opportunity for data-driven low-altitude coverage prediction. However, collecting extensive low-altitude road test data is cost-prohibitive, often yielding only sparse samples per BS. This scarcity results in two primary challenges: imbalanced feature sampling due to limited variability in high-dimensional operational parameters against the backdrop of substantial changes in low-dimensional sampling locations, and diminished generalizability stemming from insufficient data samples. To overcome these obstacles, we introduce a dual strategy comprising expert knowledge-based feature compression and disentangled representation learning. The former reduces feature space complexity by leveraging communications expertise, while the latter enhances model generalizability through the integration of propagation models and distinct subnetworks that capture and aggregate the semantic representations of latent features. Experimental evaluation confirms the efficacy of our framework, yielding a $7\%$ reduction in error compared to the best baseline algorithm. Real-network validations further attest to its reliability, achieving practical prediction accuracy with MAE errors at the 5dB5\ $\mathrm${dB} level.
@@ I Introduction
TABLE I: Difference between Terrestrial and Low-altitude Networks
Main Path Beam Impact Coverage
Low-altitude Networks Direct Path Big 3D
Terrestrial Networks Multipath Small 2D
With the burgeoning development of the low-altitude economy, researchers have redirected their attention to LANC prediction[2]. As shown in Table I, there are three main differences between terrestrial and low-altitude networks. Unlike terrestrial networks, low-altitude networks will be planned in airspace with fewer buildings, leading to the predominance of received signals from direct paths. Consequently, the actual LANC is largely influenced by the multi-beam capabilities of base stations (BSs) and the collaborative effect of multiple BSs[20]. This shifts the focus from conventional environmental scattering to base station antenna configurations and beamforming patterns. Moreover, many terrestrial models treat the vertical dimension as negligible, making them unsuitable for the three-dimensional nature of low-altitude networks. The traditional terrestrial solutions, which are primarily designed for ground-level environments, are inadequate for addressing these challenges. As a result, there is a clear need for novel approaches specifically tailored to the unique requirements of low-altitude networks.
@@
  • Experiments with Real-world Network Data: We validated the proposed framework through real-world low-altitude data collection and network training conducted in Nanchang, Jiangxi Province, China. Ablation studies demonstrate our framework reduces the mean absolute error (MAE) by approximately $20\%$ compared to traditional DNNs. Furthermore, our method outperforms the best baseline algorithm with a $7\%$ reduction in error. In practical LANC prediction of an unknown area, the framework achieves a measured MAE error at the 5dB5\ $\mathrm${dB} level, demonstrating its reliability and applicability in real-world scenarios.
@@ II-A BS and Its Operational Parameters
In scenarios such as BS pre-planning or where low-altitude sparse data collection is absent, predicting low-altitude coverage based on operational parameters becomes a prevalent approach. For such a BS of interest, the operational parameters crucially linked to coverage can be systematically categorised based on their roles in determining coverage. Given the diverse nature of BS parameters, which number in the hundreds across physical and network layers, it is essential to establish clear screening and categorisation criteria for a robust coverage prediction framework.
To address this challenge, we employ a systematic parameter selection approach guided by the principle of coverage correlation. Specifically, we retain only those parameters that directly influence the electromagnetic propagation characteristics, beam pattern, or signal strength distribution in the coverage area. This criterion enables us to exclude parameters irrelevant to physical coverage, such as core network latency, backhaul configuration, or higher-layer protocol settings, which do not affect the radio propagation environment. Through this screening process, we distil the extensive parameter space into a manageable set of coverage-critical parameters.
Following parameter selection, we establish a logical categorisation framework that organises the selected parameters according to their distinct roles in coverage determination. This categorisation yields four functional groups, each addressing a specific aspect of network coverage:
@@
For ease of description, the aforementioned categories and the included operational parameters are defined in Table II.
TABLE II: Category of Operational Parameters and Corresponding Mathematical Expression
@@ II-B Low-altitude Coverage for Area of Interest
To further explore the LANC provided by the BS, consider a nearby area with dimensions $U\times W\times V$ m3, corresponding to its longitudinal, latitudinal, and altitudinal extents. This area can be subdivided into $N$ cubes, where $N=N_{U}\times N_{W}\times N_{V}$ and $N_{U}$, $N_{W}$, and $N_{V}$ represent the counts of discretization along the longitude, latitude, and altitude axes, respectively. The set of cubes can thus be defined as $\mathbf{R}=\{1,\cdots,r_{n},\cdots,r_{N}\}$. For cube $r_{n}$, the absolute position vector can be defined as $\mathbf{x^{(R)}_{n}}=[x^{(R)}_{n,1},x^{(R)}_{n,2},x^{(R)}_{n,3}]^{T}$, where $x^{(R)}_{n,1}$, $x^{(R)}_{n,2}$ and $x^{(R)}_{n,3}$ are longitude, latitude, and altitude, respectively. Furthermore, the absolute position array for all cubes is defined as $\mathbf{X^{(R)}}=[\mathbf{x^{(R)}_{1}},\cdots,\mathbf{x^{(R)}_{n}},\cdots,\mathbf{x^{(R)}_{N}}]\in\mathbb{R}^{3\times N}$. Next, we will further elaborate on signal strength measurements and introduce cube-level coverage.
In practical implementation, signal strength measurements are obtained through low-altitude drive testing procedures using dedicated measurement terminals provided by network equipment manufacturers. These terminals systematically collect geographical coordinates, monitored frequency bands, Physical Cell Identity (PCI) of serving base stations, and comprehensive beam-specific signal measurements.
To establish the theoretical foundation for low-altitude coverage analysis, it is essential to understand the beam structure in 5G NR networks. The current 5G protocol defines three primary downlink beam categories: (i) SSB beams (static beams utilized for initial terminal access and synchronization procedures), (ii) CSI-RS beams (dynamic beams employed for channel state information estimation), and (iii) PDSCH beams (dynamic beams dedicated to data transmission). For low-altitude network coverage prediction, terminal connectivity to target base stations is predominantly determined by SSB beam quality, as these beams establish the fundamental synchronisation and access framework required for initial network attachment.
SSB-RSRP (Synchronisation Signal Block - Reference Signal Received Power) quantifies the signal strength of individual SSB beams transmitted by a base station. Within each base station’s operational framework, multiple SSB beams are transmitted using time-division multiplexing across different spatial directions to provide comprehensive coverage. SS-RSRP represents the maximum SSB-RSRP value among all transmitted SSB beams from a specific base station, effectively indicating the strongest achievable signal strength for terminal connectivity at any given spatial location.
Determined by the AAU type, the BS will emit $M$ beams. For any given cube $n$, $M$ SSB-RSRP measurements can be obtained, along with a maximum SSB-RSRP, denoted as SS-RSRP. These measurements are mathematically represented as $\mathbf{p_{n}}=[p_{n,0},p_{n,1},\cdots,p_{n,m},\cdots,p_{n,M}]^{T}\in\mathbb{R}^{(M+1)\times 1}$, where $p_{n,0}$ represents the SS-RSRP for the $n$-th cube, and $p_{n,m}$ ($m=1,2,\ldots,M$) denotes the $m$-th individual SSB-RSRP measurement for the $n$-th cube. Consequently, the comprehensive coverage characterisation for all cubes within the target area is expressed as $\mathbf{P}=[\mathbf{p_{1}},\cdots,\mathbf{p_{n}},\cdots,\mathbf{p_{N}}]\in\mathbb{R}^{(M+1)\times N}$.
@@ II-C Problem Formulation
For the whole area, the prediction of $\mathbf{P}$ can be denoted as $\mathbf{\hat{P}}=[\mathbf{\hat{p}_{1}},\cdots,\mathbf{\hat{p}_{n}},\cdots,\mathbf{\hat{p}_{N}}]\in\mathbb{R}^{(M+1)\times N}$. Finally, the problem of coverage prediction can be formulated as
[…]
@@ III-A Expert-knowledge-based Feature Compressing
where $\theta_{n}^{(H)}$ and $\theta_{n}^{(V)}$ are the horizontal and vertical angles of cube $r_{n}$ relative to the BS’s true north direction, respectively. The elements $\Delta x^{(R)}_{n,1}$, $\Delta x^{(R)}_{n,2}$, and $\Delta x^{(R)}_{n,3}$ represent the first, second, and third components of $\Delta\mathbf{x^{(R)}_{n}}$, respectively. These components signify the differences in longitude, latitude, and altitude dimensions between cube $r_{n}$ and the considered BS. Eq. (4) returns the horizontal relative angle of $r_{n}$ with respect to the true north direction of the considered BS, with a range of $\pi,\pi]$ (in radians).
@@
$D_{n}=\sqrt{{\Delta x^{(R)}_{n,1}}^{2}+{\Delta x^{(R)}_{n,2}}^{2}+{\Delta x^{(R)}_{n,3}}^{2}}.$ (8)
@@
Here, $p^{(T)}$ represents the transmission power of the SSB signal, $G^{(T)}_{n,m}$ denotes the directional gain of the BS for the $m$th beam at grid $r_{n}$, and $G^{(R)}$ is the gain of the receiver antenna. $PL_{n,m}$ is the path fading between cube $r_{n}$ and corresponding beam mm. [^3] In the context of LANC, different beams may follow distinct paths, leading to varied path fading. Given that subsequent experiments are conducted using the same single-antenna UAV, without loss of generality, we simplify the receiver antenna to an omnidirectional antenna.
Theoretically, $p^{(T)}$ can be obtained from $x^{(A)}_{1}-10\log_{10}x^{(A)}_{2}$. However, due to the actual SSB signal bandwidth utilization not being 100% and its susceptibility to the BS’s carrier frequency and bandwidth, represented by $\sigma$ [^4] In commercial 5G networks, the bandwidth utilisation parameter $\sigma$ varies according to channel bandwidth and subcarrier spacing (SCS) configurations as specified in 3GPP standards. Based on practical deployments, $\sigma$ typically ranges from approximately 0.79 to 0.98, with higher values achieved in wider bandwidth allocations. For instance, with 15 kHz SCS, $\sigma$ ranges from 0.90 (5 MHz) to 0.972 (50 MHz), while 30 kHz SCS achieves values from 0.792 (5 MHz) to 0.983 (100 MHz). In our experimental validation, we consider network configurations where $\sigma$ ranges from 0.954 to 0.983, representing high-efficiency operating conditions typical of contemporary 5G deployments. This practical grounding ensures our coverage prediction model accurately reflects real-world system performance., the actual $p^{(T)}$ is given by
@@
By incorporating expert knowledge, the transformation of the prediction task has addressed the issue of imbalanced feature variation. Simultaneously, it significantly reduces the dimensionality of the input variables, thereby ensuring an enhancement in coverage prediction performance. The specification of practical $\sigma$ ranges further ensures that our model parameters reflect realistic system operating conditions, enhancing the reliability of coverage predictions for deployment planning. Next, we will introduce the corresponding disentangled representation network design.
@@ III-B Disentangled-network-based Feature Representation
Figure 4: Full model and prior constrained model.
@@ IV-A2 Decomposing model for comparable learning capability
We try to establish that our proposed prior-constrained model structure can maintain a comparable learning capability to traditional DNNs. From an angle of deep learning practitioners, knowing how $\mathbf{x}$ is independently composed of $\mathbf{x}^{1T},\dots,\mathbf{x}^{KT}]$, we propose to design dedicated model heads for each of the components as shown in Fig. 4. Then, to make use of the full data, we aggregate the output of the heads as the final feature by concatenation. This procedure is equivalent to a constrained model. To see this, take a toy example of the linear layer in a full model, transforming a vector of $d$ dimensions into one of $d_{1}$ dimensions. The weight of this layer can be viewed as a matrix $\mathbf{w}\in\mathbb{R}^{d\times d_{1}}$. To gain a corresponding constrained model, some dimension of $\mathbf{w}$ can be constrained to $\{0\}$.
@@
$C$ is a binary random variable, takes value $1$ if a $\mathbf{w}\in\mathcal{W}_{c}$ else 0.
@@
TABLE III: Simple Description of Dataset
Area Index Longitude Latitude Length (m) Width (m) Height (m) Sample Number BS Number Size
1 115.69 28.61 4006 4273 150,300,500 43919 33 9.31MB
@@
TABLE IV: Benchmarks for Ablation Study
@@ V-A Dataset
To evaluate the performance of the proposed method, initial low-altitude road tests were conducted across three distinct areas within Nanchang City, Jiangxi Province. For these tests, a UAV equipped with a road test terminal flew at varying altitudes (150m, 300m, 500m). The terminal was capable of receiving multiple SSB-Beams’ RSRP and PCI values from the downlink of the serving cell BSs. Utilizing the PCI values, the terminals identified and matched the nearest BS with the corresponding PCI within the area. The samples collected from each region during the tests are presented in Table III.
@@ V-C1 Network Structure Comparison
To validate the advanced nature of the proposed network structure, we established three benchmarks listed in Table IV for ablation studies. In all benchmarks, the number of neurons in the hidden layers is set to 256, with the ReLU function serving as the activation function. For all benchmarks, the learning rate is set to 0.001, with a maximum of 100 epochs. Early stopping is implemented, terminating training when the loss on either the training or validation set decreases by no more than 1% over ten consecutive epochs. For benchmark 1, the hidden layer counts for distance fading, frequency fading, and antenna gain networks are all 5. In contrast, benchmarks 2 and 3 feature 6 hidden layers. Compared to the benchmark proposed in this paper, benchmark 2 utilizes decoupled features as inputs and employs an MLP to perform predictions as depicted in Eq. (11). Conversely, benchmark 3 employs original features as inputs and uses an MLP to execute predictions according to Eq. (2).
@@
Figure 6: Boxplot of MAE performance for different algorithms across 20 random splits at sampling rates of $10\%,20\%,30\%,40\%,50\%,60\%,70\%]$, totally 140 MAE data points per algorithm.
Figure 7: Boxplot of MAE performance for different algorithms across 20 random splits at sampling rates of $1\%,2\%,3\%,4\%,5\%,6\%]$, totally 120 MAE data points per algorithm.
To demonstrate the effectiveness of the disentangled feature and network modules, we initially conducted experiments with 20 different dataset splits at sampling rates of $10\%,20\%,30\%,40\%,50\%,60\%,70\%]$. For each sampling rate, we plotted the average MAE across benchmarks 1-3, as illustrated in Fig. 5. It is observable that the prediction errors for benchmarks 3, 2, and 1 progressively decrease, which validates the efficacy of the module design proposed in this study. Furthermore, as the sampling rate increases, the trends of the MAE curves for all three methods are consistently aligned, with overall performance gradually improving. This indicates that with an increase in training dataset size, the algorithms can learn a more precise distribution of the original dataset.